19.Find the capacitance of Earth. The radius of the Earth is 6,400 km By how much will the Earth's potential change if it receives a quantity of electricity equal to 1C?


1
Expert's answer
2022-01-31T07:35:03-0500

Solution

Radius of earth R=6400Km

the capacitance of Earth is given by

"C=4\pi \epsilon _0R\\=4\times3.14\times8.85\times10^{-12}\times 6400000"

C=0.71mF

Now

Earth's potential change 

"V=\frac{C}{Q}\\=\frac{0.71*10^{-3}}{1}\\=0.71mV"